1. Suppose we are selling hot dogs during a basketball game. A hot dog sells for $15 each but only costs $5 each to make. If we run out of hot dogs during the game, it will be impossible to get more. On the other hand, a leftover hot dog has a value of S1 for each hot dog. Assume that we believe the fans would buy 100 hot dogs with probability 3/10, 110 hot dogs with probability 2/10, 120 hot dogs with probability 2/10, 130 hot dogs with probability 1/10, and 140 hot dogs with probability 2/10. (a) Find the expected demand. (b) If 120 hot dogs are prepared, what is the expected profit? (c) What is the corresponding expected (potential) cost?
Answer #1

Solution :-

Given data :

A hot dog sells for $15 each but only costs $5 each to make.

A leftover hot dog has a value of $1 for each hot dog.

demand and probabilities are

1. x_1 = 100 hot dogs with probability P(x_1) = 3/10,

2. x_2 = 110 hot dogs with probability P(x_2) = 2/10,

3. x_3 = 120 hot dogs with probability P(x_3) = 2/10,

4. 04 = 130 hot dogs with probability P(04) = 1/10,

5. x_5 = 140 hot dogs with probability P(x_5) = 2/10.

( a ) :-

Here we need to calculate the expected demand.

i.e, E[X].

E[X] = sum x_i*P(x_i)

EX| = 100% (3/10) +110% (2/10) +120 * (2/10) +130*(1/10) +140(2/10)

EXI = 30 22 24 13 28

EX 17

herefore Expected demand, EX 17

( b ) :-

Here we need to calculate the expected profit, If 120 hot dogs are prepared.

When a hot dog sold, will earn a profit of ( $15 - $5 ) = $10

When a hot dog unsold, will be a loss of ( $5 - $1 ) = $4

The demand is more that too will be a potential loss, If the hot dogs are finished and this will be a loss of potential profit of $10.

If the 120 hot dogs are prepared then, the profit will be also have the probability distribution according to the distribution of demand.

Consider the demand is 100 , then 20 hot dogs will be left.

and the profit = (100*10)-(20*4)

profit = 1000-80

   profit = 920

Consider the demand is 110 , then 10 hot dogs will be left.

and the profit = (110*10)-(10*4)

profit = 1100-40

profit = 1060

Then the expected profit becomes as

Expected profit =920 * (3/10) +1060 *(2/10) +1200 (2/10) +1200* (1/10) 1200*(2/10)

Expected profit = 920 * 0.3+ 1060 0.21200 0.2 -1200 0.11200 * 0.2

Expected profit = 276+212+240+120+240

Expected profit = $1088

herefore  Expected profit = $1088

( c ) :-

Here we need to calculate  the corresponding expected (potential) cost.

Here we know the formula,

Expected (potential) cost = Expected demand * Cost

Expected (potential) cost = 117 * $5

Expected (potential) cost = $585   

herefore Expected (potential) cost = $585